OpenStudy (anonymous):

PLEASE HELP What is the standard form of the expression i/2+i? (i divided by 2+i)

hartnn (hartnn):

multiply and divide by conjugate, that is 2-i

OpenStudy (anonymous):

thanks again

hartnn (hartnn):

u got it ?

OpenStudy (anonymous):

actually im having trouble multiplying it out

hartnn (hartnn):

i(2-i) is numerator = ? (2+i)(2-i) = 4-i^2 = .. ?

hartnn (hartnn):

u know i^2= .... ?

OpenStudy (anonymous):

so far I have 2i-i^2 over 4-i^2

OpenStudy (anonymous):

i got 1/2i+1

hartnn (hartnn):

huh ? i^2 = -1 u knew this ?

OpenStudy (anonymous):

yea i know that

hartnn (hartnn):

4-i^2 = 4-(-1) = 5

OpenStudy (anonymous):

wait so what would the answer be, kinda confused

OpenStudy (anonymous):

so is it 2/5i+1/5?

hartnn (hartnn):

i(2-i)/5 = 2i/5 - i^2/5 = 2i/5 + 1/5

hartnn (hartnn):

yes, thats correct :)

OpenStudy (anonymous):

oh ok, thanks a lot, better teacher than my own at school lol

hartnn (hartnn):

welcome again :)